OpenStudy (anonymous):

calculate the density of a solid with the following data: Volume of water: 48.4 ml Volume of water and solid: 62.7 ml Mass of solid: 123.4g Please Help, Thank You :) calculate the density of a solid with the following data: Volume of water: 48.4 ml Volume of water and solid: 62.7 ml Mass of solid: 123.4g Please Help, Thank You :) @Physics

OpenStudy (beefarm):

volume of the solid = (volume of water and solid)-(volume of water) = 14.3 density = m/v = 123.4/14.3 g/ml Also, you'll probably want to convert this into SI, units. (kg/cm^3, I think)